If MARR = 8%, find the following for cash flow below:

EOY

Net CF

0

-25,000

1

5,000

2

5,000 + X

3

5,000 + X

4

5,000 + X

5

5,000 + X

6

5,000 + X

7

5,000 + X

a) find X if the AW = 6000

b) Find X if the ERR is 12% (you may assume X is positive)

c) Find X if the Benefit/Cost Ratio is 1.12 (you may assume X is positive)
Answer #1

A.

B. ERR = 12%

55,267.035 = 44,614.02 + 7.3359X

7.3359X = 10,653.02

X = $ 1,452 (Approximately)

C. Benefit cost ratio = 1.12

1.12×25000 = 5000 × 5.20637+ X × 4.622879 ×0.9259

28,000 = 26,031.85 + 4.28044X

4.28044X = 1,968.15

X = $ 460 (Approximately)

Please contact if having any query will be obliged to you for your generous support. Your help mean a lot to me please help. Thank you.